OpenStudy (wach):

Emma took a certain 100-point test twice. The points received on the test can be converted to a rating from according to the formula, R=1.4P+23, where R represents the rating and P represents the points received. How many more points did Emma receive on her second test compared to her first if her rating increased by 63? I do not need an answer, the correct answer is 45. I would like a formula to calculate this along with somewhat of an explanation.

OpenStudy (anonymous):

my name is emma :)

OpenStudy (ganpat):

R=1.4P+23 R2 - R1 = 63.. So, (1.4P2 + 23 ) - (1.4P1 + 23) = 63 1.4P2 -1.4P1 = 63 1.4 (P2 - P1 ) = 63 P2 -P1 = 45...

OpenStudy (ganpat):

@Emmaleee : You have improved a lot :-p :D...

OpenStudy (anonymous):

how did i improve ? in what ?

OpenStudy (wach):

Thanks so much, Ganpat! I was afraid that no one would help me. Good explanation!

OpenStudy (ganpat):

lol !! Your first comment, " My name is Emma " :D Just Kidding,

OpenStudy (ganpat):

@wach : no worries dude, there are many experts around.. m just a small fish..